OpenStudy (anonymous):

please help! How would I find x in the proportion 4x-3/x = 2x/x? I'm confused because there are x's in all the denominators and numerators

OpenStudy (anonymous):

does it mean (4x-3)/x=2x/x

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\frac{ 4x-3 }{ x } =\frac{ 2x }{ x }\]

OpenStudy (anonymous):

(4x-3)/x=2x/x multiply both sides by x to get ride of the fractions 4x-3=2x -4x -4x -3=-2x divide both sides by -2 3/2=x

OpenStudy (anonymous):

Oh okay, I understand. Thank you so much for your help! I appreciate it!

OpenStudy (anonymous):

np